If X is normally distributed with mean 15 and variance 16, find mean & standard deviation of Y = 3X +ย 3


1
Expert's answer
2021-11-18T12:23:12-0500

"E(Y)=E(3X+3)=3E(X)+3=3*45+3=138."

"Var(Y)=Var(3X+3)=3^2Var(X)=9*16=144."

"\sigma(Y)=\sqrt{Var(Y)}=\sqrt{144}=12."